OpenStudy (kj4uts):

1. Consider the equation cos(2.13t)=0.5. Find the smallest positive solution in radians and round your answer to four decimal places. 2. Consider the equation tan(2.81t)=10.0. Find the smallest positive solution in radians and round your answer to four decimal places.

OpenStudy (kj4uts):

I think this is how you solve the problems? 1. cos^-1(0.5)/2.13 = 0.4916 2. tan^-1(10)/2.81 = 0.5235

OpenStudy (loser66):

cos (? )=0.5?

OpenStudy (kj4uts):

What do you mean?

OpenStudy (loser66):

on unit circle, which angle gives you cos (that angle ) = 0.5?

OpenStudy (loser66):

don't know?

OpenStudy (kj4uts):

No, the way I solved it was arccos(0.5)/2.13 = 0.4916

OpenStudy (loser66):

ok, it's up to you.

OpenStudy (kj4uts):

Does that look right?

OpenStudy (kj4uts):

What was the way you were going to do it?

OpenStudy (loser66):

Yes, you are correct

OpenStudy (loser66):

My way: cos (pi/3) = 0.5 = cos (2.13 t) Hence pi/3 = 2.13 t that gives \(t = (\pi/3)/2.13 = 0.4916\)

OpenStudy (kj4uts):

oh ok the pi/3 is the angle on unit circle?

OpenStudy (kj4uts):

would you do the same for number 2 with tan?

OpenStudy (loser66):

Nope, for tan, you need your method.

OpenStudy (kj4uts):

Does my tan answer look right?

OpenStudy (kj4uts):

Ok thank you very much for your time and help :)

OpenStudy (loser66):

It shows t = 29.996

OpenStudy (kj4uts):

is that in radians or degrees?

OpenStudy (loser66):

degree

OpenStudy (kj4uts):

ok I think 0.5235 is the answer in radians

OpenStudy (kj4uts):

0.5235 was correct, thank you for your time and help :)